# SpotFormer: Multi-Scale Spatio-Temporal Transformer for Facial Expression Spotting

### v1

@@ 3.4 Post-processing
Once the series of output probabilities $P=(p_{i})_{i=1}^{N}$ for the entire video $V$ are obtained, we perform MaE spotting and ME spotting independently, following the methodology outlined in [23]. We will elaborate on the ME spotting process, with the MaE spotting process following a similar method. Initially, we spot all potential ME apex frames $U^{\mathrm{mi},\mathrm{apex}}$ based on the criterion $p_{l}^{\mathrm{mi},\mathrm{apex}}>\theta_{\mathrm{apex}}$, where $\theta_{\mathrm{apex}}$ is a threshold. For each spotted apex frame $u^{\mathrm{mi},\mathrm{apex}}_{l}\in U^{\mathrm{mi},\mathrm{apex}}$, we select the onset frame with the highest onset probability from the left side of the apex frame within the range of $[l-\frac{k^{\mathrm{mi}}}{2},l-\frac{j^{\mathrm{mi}}}{2}]$. Similarly, we select the offset frame with the highest offset probability from the right side of the apex frame within the range of $[l+\frac{j^{\mathrm{mi}}}{2},l+\frac{k^{\mathrm{mi}}}{2}]$, where $k^{\mathrm{mi}}$ and $j^{\mathrm{mi}}$ represent the average duration and minimum duration of a ME, respectively. As a result, a ME proposal $\phi_{l}$ is obtained, including the onset frame, offset frame, and expression type. Subsequently, we assign a score $s_{l}=p_{b}^{\mathrm{mi},\mathrm{onset}}\times p_{l}^{\mathrm{mi},\mathrm{apex}}\times p_{d}^{\mathrm{mi},\mathrm{offset}}$ to $\phi_{l}$, where $b$ and $d$ denote the frame indices of the onset frame and offset frame selected by the aforementioned rule, respectively.
@@ 4 Experiments
In this section, we first introduce the experimental setup (Section 4.1) used in the paper. We then compare our method with state-of-the-art methods to demonstrate the effectiveness of SpotFormer (Section 4.2). Next, we thoroughly evaluate our proposed modules in SpotFormer in the ablation study (Section 4.3). We then investigate the effects from the different designs in SpotFormer (Section 4.4). Finally, we present the detailed expression spotting results and a discussion of the results (Section 4.5).
@@ 4.1 Experimental setup
Datasets. We conduct experiments on two datasets: SAMM-LV [61] and CAS(ME)2 [62]. The SAMM-LV dataset includes 147 raw long videos with 343 MaE clips and 159 ME clips. The CAS(ME)2 dataset includes 87 raw long videos with 300 MaE clips and 57 ME clips. Since the frame rate of the SAMM-LV dataset is 200fps while the frame rate of the CAS(ME)2 datasets is 30fps, we subsample every 7th frame from the SAMM dataset to align the frame rates of both datasets.
Evaluation metric. Following the protocol of MEGC2021, we employ a leave-one-subject-out cross-validation strategy in our experiments. The true positive (TP) is defined based on the Intersection over Union (IoU) between the proposal and the ground-truth expression clip. Specifically, given a ground-truth expression clip and its expression type, we compare it with all expression proposals that have the same estimated expression type. An expression proposal $W_{\mathrm{Proposal}}$, is considered a TP when it satisfies the following condition:
@@
Training details. We use the AdamW optimizer [63] to optimize our model, setting the learning rate to 0.0002, $\beta_{1}$ to 0.7, and $\beta_{2}$ to 0.9. The temperature parameter $\tau$ in (4) is set to 0.5. We train our model for 100 epochs with a batch size of 512.
@@ 4.2 Comparison with state-of-the-art methods
We first compare our method with the state-of-the-art methods on the SAMM-LV and CAS(ME)2 datasets, and the results are shown in Table I. We report the F1-score for MaE spotting, ME spotting, and overall performance. For the overall performance, our method achieves F1-scores of 0.4401 on the SAMM-LV dataset and 0.4841 on the CAS(ME)2 dataset, which outperforms other state-of-the-art methods by 13.4% and 6.2%, respectively. For the MaE spotting, our method achieves an improvement of 3.6% on the SAMM-LV dataset compared to other methods. It is important to emphasize our method’s remarkable effectiveness in ME spotting. The results demonstrate a substantial enhancement with an 49.4% improvement on the SAMM-LV dataset and a 7.8% improvement on the CAS(ME)2 dataset compared to other state-of-the-art methods. The results show our method’s ability to capture subtle motions that exist in MEs and alleviate the impact of irrelevant motions.
@@ 4.3 Ablation studies
We conduct ablation studies to evaluate our proposed modules and Table II shows the experimental results. The acronyms SW-MRO and FLGP denote specific modules in our model. Specifically, SW-MRO involves calculating sliding window-based multi-resolution optical flow features instead of optical flow between adjacent frames, while FLGP integrates facial local graph pooling operation into SpotFormer for multi-scale spatial feature learning instead of global average pooling over facial graphs. The results show the effectiveness of each proposed module. Notably, the introduction of the SW-MRO enhances the extracted motion features, particularly in revealing subtle motions that exist in MEs. This enhancement leads to a significant overall performance improvement of 17.0%/17.7% on the SAMM-LV and CAS(ME)2 datasets. Especially in ME recognition, it has led to an improvement of 69.2%/200.1% on the SAMM-LV and CAS(ME)2 datasets. The introduction of facial local graph pooling operation for multi-scale spatial feature learning results in a further improvement of 3.0%/3.4% on the SAMM-LV and CAS(ME)2 datasets compared to using global average pooling. This demonstrates the superior representational capabilities of our proposed model.
TABLE II: Results of the ablation study on the effectiveness of the proposed modules.
@@ 4.4.1 Multiple model architectures
To verify the effectiveness of the SpotFormer, we created multiple model architectures that consider multi-scale spatio-temporal feature learning differently as variants of SpotFormer and then conducted a comprehensive comparison among them. Here, the original architecture of SpotFormer is called Simul-ST for clarity since it simultaneously models spatio-temporal relations and enables the comprehensive capture of both spatial and temporal information. The architectures of variants are described as follows:
Figure 6: Overview of the variants of the proposed model architectures. (a) Seq-ST encodes each graph and then models temporal variations; (b) Paral-ST learns temporal variations in various graph scales and performs multi-scale feature fusion; (c) Simul-TS learns temporal variations first and then learns spatial relations simultaneously.
Seq-ST. This model sequentially learns spatial features first, followed by temporal features. As shown in Fig. 6 (a), we first employ two spatial graph attention layers, each followed by an FLGP layer. After downsampling the facial graph structure to one node, we add temporal node attention layers, where each is followed by a temporal downsampling layer until temporal resolution is downsampled to 1. Seq-ST is similar to a general TAL architecture, which encodes spatial information into a feature vector and then learns temporal information to detect potential action instances. In both TAL and FES, temporal information is vital for detecting action instances; such a model can initially focus on spatial details, capturing fine-grained information, and then capturing temporal dynamics.
Paral-ST. Considering that Seq-ST lacks the consideration of temporal information in lower graph scales, Paral-ST parallelizes temporal learning across different spatial scales and fuses them. Specifically, as described in Fig. 6 (b), we perform temporal node attention and temporal downsampling operations at each facial graph scale. At each temporal scale, we fuse temporal features from different graph scales by incorporating the output of the temporal node attention layer into the higher graph scale. The incorporation is achieved through a CNN layer followed by an element-wise addition operation. In Paral-ST the independent extraction of temporal dynamics for each spatial graph scale and the multi-scale feature fusion ensure a comprehensive understanding of temporal dynamics.
TABLE III: Comparison of multiple model architectures on SAMM-LV and CAS(ME)2.
@@
Simul-TS. Would it be better to prioritize temporal node attention first? With such consideration, Simul-TS begins with a focus on temporal dynamics and gradually introduces spatial information. As described in Fig. 6 (c), in each attention block, we first perform temporal node attention, followed by spatial graph attention. Subsequently, we perform temporal downsampling and spatial downsampling to generate high-level features. By learning spatial information later in the process, Simul-TS treats spatial information as more important.
TABLE IV: Comparison of different temporal window sizes (receptive fields) on SAMM-LV.
@@
We validate and compare our proposed multiple multi-scale spatio-temporal model architectures on the SAMM and CAS(ME)2 datasets, and the experimental results are shown in Table III. With the exception of Paral-ST, our models use the same number of parameters for a fair comparison.
TABLE V: Results of the ablation study on supervised contrastive learning and the choice of hyperparameter.
@@
4.4.2 Temporal window size
We further explore how much temporal information is needed to accurately evaluate one frame. Specifically, we compare different temporal window sizes (i.e., temporal receptive field) for the SW-MRO on the SAMM-LV dataset, and the experimental results are presented in Table IV. The results indicate that the optimal performance is obtained when the temporal window size is set to 17, which demonstrates the effectiveness of our core idea: perceiving complete MEs and distinguishing between MaEs and MEs. Specifically, the temporal boundary for distinguishing MaEs and MEs is 0.5 seconds, which corresponds to 15 frames when the frame rate is set to 30fps. A receptive field of 17 frames proves sufficient to perceive a complete ME and amplify the motion information that exists in MEs. When the model can accurately spot MEs, windows that exhibit greater intensity of motion, or windows that exhibit a continuous expression and are hard to perceive a trend of this expression occurring or ceasing within a short period, are more likely to contain MaEs.
Based on this idea, the model can capture the temporal variations necessary to distinguish between general MaEs and MEs. A smaller window size cannot provide sufficient temporal information, while an increase in the window size results in more frames, potentially leading to information redundancy and increased ambiguity in distinguishing between MaEs and MEs. Moreover, larger temporal window sizes may cause more severe head movement problems, negatively impacting overall performance.
@@ 4.4.2 Temporal window size
Figure 8: Qualitative results of our method on SAMM-LV. The example video is “$006\_1$”, whose frame rate has been downsampled to 30fps.
TABLE VI: Detailed spotting results of the proposed method on SAMM-LV and CAS(ME)2.
@@ 4.4.3 Supervised contrastive loss and hyperparameter $\lambda$
Table V shows the experimental results of introducing supervised contrastive learning into our model and the impact of the hyperparameter $\lambda$ in (5), which is set to balance classification and supervised contrastive learning. We observe that the introduction of the supervised contrastive loss enables our model to better recognize the boundaries distinguishing between different types of expressions, resulting in an improvement of 6.0%/16.3% on the SAMM-LV and CAS(ME)2 datasets. The optimal value for $\lambda$ is 0.005, increasing $\lambda$ beyond this value starts to impact standard classification, leading to a decrease in spotting accuracy.
Fig. 7 shows the qualitative comparison between scenarios with and without supervised contrastive learning. In the inference phase, we randomly sampled frames and applied principal component analysis (PCA) [64] to analyze the distribution of various expression classes. Subsequently, we marked each frame with its corresponding ground-truth expression label. In Fig. 7 (a) and (b), we examine the PCA distribution of specific MaE frames and ME frames with and without supervised contrastive learning. Moving to Fig. 7 (c) and (d), we compare the PCA distribution of specific neutral frames and ME frames with and without supervised contrastive learning. The results indicate that, in the absence of supervised contrastive learning, our model may face challenges distinguishing certain MaE frames from ME frames due to subtle differences. This leads to mixed distributions and misclassification. Additionally, neutral frames may be mistakenly classified as ME frames, as optical flow noise introduces slight differences that the model treats as ME motion information, resulting in incorrect ME proposals detection. With the introduction of supervised contrastive learning into our model, the domain discrepancy increases, thus significantly improving the accuracy of expression recognition.
@@ 4.5 Detailed results
Table VI shows the detailed results on the SAMM-LV and CAS(ME)2 datasets. Our method achieves a recall rate of approximately 0.5 in MaE spotting and 0.4 in ME spotting. The exception is for ME spotting on the CAS(ME)2 dataset. The results indicate that it achieves a high precision score while a low recall score, which may be due to several factors. One primary reason is the scarcity of ME data and data imbalance in the CAS(ME)2 dataset, which only contains 57 ME clips, some of which involve only eye blinking—considered irrelevant motions for facial expression spotting. Furthermore, owing to our model’s strong discriminatory capability and the incorporation of supervised contrastive learning, the model cautiously evaluates ME frames based on limited data. Consequently, it achieves a high precision rate but a low recall rate.

### v2

@@ 3.4 Post-processing
Once the series of output probabilities $P=(p_{i})_{i=1}^{N}$ for the entire video $V$ are obtained, we perform MaE spotting and ME spotting independently, following the methodology outlined in [leng2022abpn]. We will elaborate on the ME spotting process, with the MaE spotting process following a similar method. Initially, we identify all potential ME apex frames $U^{\mathrm{me},\mathrm{apex}}$ based on the criterion $p_{l}^{\mathrm{me},\mathrm{apex}}>\theta_{\mathrm{apex}}$, where $\theta_{\mathrm{apex}}$ is a threshold. For each spotted apex frame $u^{\mathrm{me},\mathrm{apex}}_{l}\in U^{\mathrm{me},\mathrm{apex}}$ with the frame index $l$, we select the onset frame with the highest onset probability from the left side of the apex frame within the range of $l-\frac{k^{\mathrm{me}}}{2},l-\frac{j^{\mathrm{me}}}{2}]$. Similarly, we select the offset frame with the highest offset probability from the right side of the apex frame within the range of $l+\frac{j^{\mathrm{me}}}{2},l+\frac{k^{\mathrm{me}}}{2}]$, where $k^{\mathrm{me}}$ and $j^{\mathrm{me}}$ represent the average duration and minimum duration of a ME, respectively. As a result, a ME proposal $\phi_{l}$ is obtained, including the onset frame, offset frame, and expression type. Subsequently, we assign a score $s_{l}=p_{b}^{\mathrm{me},\mathrm{onset}}\times p_{l}^{\mathrm{me},\mathrm{apex}}\times p_{d}^{\mathrm{me},\mathrm{offset}}$ to $\phi_{l}$, where $b$ and $d$ denote the frame indices of the onset frame and offset frame selected by the aforementioned rule, respectively.
@@ 4 Experiments
In this section, we first introduce the experimental setup (Section 4.1) used in the paper. We then compare our method with state-of-the-art methods to demonstrate the effectiveness of SpotFormer (Section 4.2). Next, we thoroughly evaluate our proposed modules in SpotFormer in the ablation study (Section 4.3). We then explore several model variants and compare them with SpotFormer to investigate its effectiveness. (Section 4.4). Finally, we present the detailed expression spotting results and a discussion of the results (Section 4.5).
@@ 4.1 Experimental setup
Datasets. We conduct experiments on two datasets: SAMM-LV [yap2020samm], CAS(ME)2 [qu2017cas], and CAS(ME)3 [9774929]. The SAMM-LV dataset includes 147 raw long videos with 343 MaE clips and 159 ME clips. The CAS(ME)2 dataset includes 87 raw long videos with 300 MaE clips and 57 ME clips. The large-scale dataset CAS(ME)3 includes 956 raw long videos with 3490 MaE clips and 1109 ME clips. Since the frame rate of the SAMM-LV dataset is 200fps while the frame rate of the CAS(ME)2 and CAS(ME)3 datasets is 30fps, we subsample every 7th frame from the SAMM dataset to align the frame rates of all three datasets.
Evaluation metric. Following the protocol of MEGC2021 [li2021fme], we evaluate our model on each dataset separately and employ a leave-one-subject-out cross-validation strategy in our experiments. The true positive (TP) is defined based on the Intersection over Union (IoU) between the proposal and the ground-truth expression clip. Specifically, given a ground-truth expression clip and its expression type, we compare it with all expression proposals that have the same estimated expression type. An expression proposal $W_{\mathrm{Proposal}}$, is considered a TP when it satisfies the following condition:
@@
Training details. We use the AdamW optimizer [loshchilov2017decoupled] to optimize our model, setting the learning rate to 0.0002, $\beta_{1}$ to 0.7, and $\beta_{2}$ to 0.9. The temperature parameter $\tau$ in Eq. (4) is set to 0.5. We train our model for 100 epochs with a batch size of 512. The facial images input for the SW-MRO feature extraction are resized to $(320,320)$. The general expression durations, $k^{\mathrm{mae}}$ and $k^{\mathrm{me}}$, are set to 60 and 16, respectively, while the minimum expression durations, $j^{\mathrm{mae}}$ and $j^{\mathrm{me}}$, are set to 16 and 4. The hidden dimension and the number of heads in the MHSA mechanism is set to 256 and 4, respectively.
@@ 4.2 Comparison with state-of-the-art methods
We first compare our method with the state-of-the-art (SOTA) methods on the SAMM-LV and CAS(ME)2 datasets, and the results are shown in Table I. We report the F1-score for MaE spotting, ME spotting, and overall performance. For the overall performance, our method achieves F1-scores of 0.4401 on the SAMM-LV dataset and 0.4841 on the CAS(ME)2 dataset, which outperforms other SOTA methods by 13.4% and 6.2%, respectively. For the MaE spotting, our method achieves an improvement of 3.6% on the SAMM-LV dataset compared to other methods. It is important to emphasize our method’s remarkable effectiveness in ME spotting. The results demonstrate a substantial enhancement with a 49.4% improvement on the SAMM-LV dataset and a 7.8% improvement on the CAS(ME)2 dataset compared to other SOTA methods.
The remarkable improvements in ME spotting are mainly attributed to our proposed SW-MRO feature. Previous methods generally extract optical flow between adjacent frames and employ either a pretrained I3D model or simple neural networks to model these coarse motion patterns. In contrast, our proposed SW-MRO feature effectively extracts subtle motions while avoiding the optical flow being dominated by head movements. Moreover, the carefully designed SpotFormer can comprehensively analyze the latent motion patterns within the SW-MRO feature. In addition, the proposed supervised contrastive learning further enhances discriminative feature learning between different expression types, leading to improved model performance.
We further validate our method on the CAS(ME)3 dataset [9774929] and the results are shown in Table II. Following LGSNet [yu2023lgsnet], we pre-process the dataset by considering only ground-truth MEs with a duration of less than 15 frames and ground-truth MaEs with a duration between 16 and 120 frames for training and validation, ensuring a fair comparison. This data pre-processing method results in 2231 ground-truth MaEs and 285 ground-truth MEs. The results demonstrate that our method outperforms LGSNet [yu2023lgsnet], especially in ME spotting, highlighting the strong generalization ability of our method.
TABLE II: Comparison with the state-of-the-art method on CAS(ME)3 in terms of F1-score.
Methods CAS(ME)3
MaE ME Overall
LGSNet [yu2023lgsnet] - 0.0990 0.2350
Ours 0.2664 0.2037 0.2559
@@ 4.3 Ablation studies
We conduct ablation studies to evaluate our proposed modules and Table III shows the experimental results. The acronyms SW-MRO and FLGP denote specific modules in our model. Specifically, SW-MRO involves calculating sliding window-based multi-temporal-resolutionoptical flow features instead of optical flow between adjacent frames, while FLGP integrates facial local graph pooling operation into SpotFormer for multi-scale spatial feature learning instead of global average pooling over facial graphs. The results show the effectiveness of each proposed module. Notably, the introduction of the SW-MRO enhances the extracted motion features, particularly in revealing subtle motions that exist in MEs. This enhancement leads to a significant overall performance improvement of 17.0%/17.7% on the SAMM-LV and CAS(ME)2 datasets. Especially in ME spotting, it has led to an improvement of 69.2%/200.1% on the SAMM-LV and CAS(ME)2 datasets. The introduction of the FLGP operation for multi-scale spatial feature learning results in a further improvement of 3.0%/3.4% on the SAMM-LV and CAS(ME)2 datasets compared to using global average pooling. This demonstrates the superior representational capabilities of our proposed model.
TABLE III: Results of the ablation study on the effectiveness of the proposed modules.
@@ 4.4.1 Multiple model architectures
To verify the effectiveness of SpotFormer, we explored multiple model architectures that consider multi-scale spatio-temporal feature learning differently as variants of SpotFormer and then conducted a comprehensive comparison among them. Here, the original architecture of SpotFormer is called Simul-ST for clarity since it simultaneously models spatio-temporal relations and enables the comprehensive capture of both spatial and temporal information. The architectures of variants are described as follows:
Figure 6: Overview of the variants of the proposed model architectures. (a) Seq-ST first employ spatial graph attention layers and FLGP operations to encode the graph of each frame to a single node, and then models temporal variations across frames; (b) Paral-ST differs from Seq-ST by modeling temporal variations at each graph scale and performing multi-scale feature fusion; (c) Simul-TS differs from SpotFormer by modeling temporal variations first and subsequently learning spatial relationships within each spatio-temporal attention block.
Seq-ST. Seq-ST sequentially learns spatial features first, followed by temporal features. As shown in Fig. 6 (a), we first employ two spatial graph attention layers, each followed by an FLGP operation for spatial downsampling. After reducing the facial graph structure to one node, we start adding temporal node attention layers, where each is followed by a temporal downsampling layer until temporal dimension is downsampled to 1. Seq-ST is similar to a general TAL architecture, which encodes spatial information into a feature vector and then learns temporal information to detect potential action instances. In both TAL and FES, temporal information is vital for detecting action instances; such a model can initially focus on spatial details, capturing fine-grained information, and then capturing temporal dynamics.
Paral-ST. Considering that Seq-ST lacks the consideration of temporal information in lower graph scales, Paral-ST parallelizes temporal learning at each spatial scale and fuses them. Specifically, as described in Fig. 6 (b), we perform temporal node attention and temporal downsampling at each facial graph scale. At each temporal scale, we fuse temporal features from different graph scales by incorporating the output of the temporal node attention layer into the higher graph scale. The incorporation is achieved through a CNN layer followed by an element-wise addition operation. In Paral-ST, the independent extraction of temporal dynamics for each spatial graph scale and the multi-scale feature fusion ensure a comprehensive understanding of temporal dynamics.
TABLE IV: Comparison of multiple model architectures on SAMM-LV and CAS(ME)2.
@@
Simul-TS. Would it be better to prioritize temporal node attention first? With such consideration, Simul-TS begins by focusing on learning temporal dynamics and gradually introduces spatial feature learning. As described in Fig. 6 (c), in each spatio-temporal attention block, we first perform temporal node attention, followed by spatial graph attention. Subsequently, we perform temporal downsampling and spatial downsampling to generate high-level features. By learning spatial information later in the process, Simul-TS treats spatial information as more important.
We validate and compare our proposed multiple multi-scale spatio-temporal model architectures on the SAMM and CAS(ME)2 datasets, and the experimental results are shown in Table IV. With the exception of Paral-ST, our models use the same number of parameters for a fair comparison.
4.4.2 Temporal window size
We further explore how much temporal information is needed to accurately evaluate one frame. Specifically, we compare different temporal window sizes (i.e., temporal receptive field) for the SW-MRO on the SAMM-LV dataset, and the experimental results are presented in Table V. The results indicate that the optimal performance is obtained when the temporal window size is set to 17, which demonstrates the effectiveness of our core idea: perceiving complete MEs and distinguishing between MaEs and MEs. Specifically, the temporal boundary for distinguishing MaEs and MEs is 0.5 seconds, which corresponds to 15 frames when the frame rate is set to 30 fps. A receptive field of 17 frames proves sufficient to perceive a complete ME and amplify the motion information that exists in MEs. When the model can accurately spot MEs, windows that exhibit greater intensity of motion, or windows that exhibit a continuous expression and are hard to perceive a trend of this expression occurring or ceasing within a short period, are more likely to contain MaEs.
Based on this idea, the model can capture the temporal variations necessary to distinguish between general MaEs and MEs. A smaller window size cannot provide sufficient temporal information, while an increase in the window size results in more frames, potentially leading to information redundancy and increased ambiguity in distinguishing between MaEs and MEs. Moreover, larger temporal window sizes may cause the optical flow to be severely dominated by head movements, negatively impacting overall performance.
TABLE V: Comparison of different temporal window sizes (receptive fields) on SAMM-LV.
@@ 4.4.2 Temporal window size
TABLE VI: Results of the ablation study on supervised contrastive learning and the choice of hyperparameter.
@@
Figure 8: Qualitative results of our method on SAMM-LV. The example video is “006_1006\_1”, whose frame rate has been downsampled to 30fps.
TABLE VII: Detailed spotting results of the proposed method on SAMM-LV and CAS(ME)2.
@@ 4.4.3 Supervised contrastive loss and hyperparameter $\lambda$
Table VI shows the experimental results of introducing supervised contrastive learning into our model and the impact of the hyperparameter $\lambda$ in Eq. (5), which is set to balance classification and supervised contrastive learning. We observe that the introduction of the supervised contrastive loss enables our model to better recognize the classification boundaries between different types of expressions, resulting in an improvement of 6.0%/16.3% on the SAMM-LV and CAS(ME)2 datasets. The optimal value for $\lambda$ is 0.005, increasing $\lambda$ beyond this value starts to impact standard classification, leading to a decrease in spotting accuracy.
Fig. 7 shows the qualitative comparison between scenarios with and without supervised contrastive learning. In the inference phase, we randomly sampled frames and applied principal component analysis (PCA) [bro2014principal] to analyze the distribution of various expression classes. Subsequently, we marked each frame with its corresponding ground-truth expression label. In Fig. 7 (a) and (b), we examine the PCA distribution of specific MaE frames and ME frames with and without supervised contrastive learning. Moving to Fig. 7 (c) and (d), we compare the PCA distribution of specific neutral frames and ME frames with and without supervised contrastive learning. The results indicate that, in the absence of supervised contrastive learning, our model may face challenges distinguishing certain MaE frames from ME frames due to subtle differences. This leads to mixed distributions and misclassification. Additionally, neutral frames may be mistakenly classified as ME frames, as the noise existing in the optical flow introduces slight fluctuations that the model may treat as ME motion information, resulting in incorrect ME proposal detection. With the introduction of supervised contrastive learning into our model, the domain discrepancy increases, thus significantly improving the accuracy of expression spotting.
@@ 4.5 Detailed results
Table VII shows the detailed results on the SAMM-LV and CAS(ME)2 datasets. Our method achieves an overall precision rate of 0.4341/0.4781 and an overall recall rate of 0.4462/0.4902 on SAMM and CAS(ME)2, respectively. Specifically, our method achieves a recall rate of approximately 0.5 in MaE spotting and 0.4 in ME spotting. The exception is for ME spotting on the CAS(ME)2 dataset. The results indicate that it achieves a high precision rate while a low recall rate, which may be due to several factors. One primary reason is the scarcity and imbalance of ME data in the CAS(ME)2 dataset, which contains only 57 ME clips. Some ME samples in CAS(ME)2 involve only eye blinking—a type of motion generally considered irrelevant for facial expression spotting. Furthermore, owing to our model’s strong discriminative capability and the incorporation of supervised contrastive learning, it carefully evaluates ME frames based on limited data. Consequently, it achieves a high precision rate but low recall. In addition, on both datasets, the precision is lower than the recall in MaE spotting, while the opposite is observed in ME spotting. This is mainly due to data imbalance and the specific characteristics of different expression classes in the FES datasets: MaEs are more frequent and generally exhibit higher intensity than MEs. This leads the model to focus more on MaEs and generate higher output probabilities for MaE spotting. As a result, more MaEs—including both TPs and FPs—are detected. In contrast, the model generates lower probabilities for ME spotting, resulting in fewer expression proposals but also fewer FPs.
It can also be observed that the overall performance on the SAMM dataset is lower than that on the CAS(ME)2 dataset. The reasons may lie in two main aspects. First, CAS(ME)2 contains very few ME samples, so the ME spotting performance contributes little to the overall performance. This explains why, even though the ME spotting performance of CAS(ME)2 is lower, the overall performance is still maintained at a high level. During both training and inference, the model tends to focus more on MaE spotting, which is relatively easier than ME spotting and has a greater impact on the overall performance. In contrast, for SAMM, since the ground-truth ME samples account for almost one-third of the dataset, ME spotting performance becomes important for the overall performance. Therefore, compared to CAS(ME)2, the model needs to pay more attention to ME spotting on SAMM, which is more challenging in the FES task. Second, the durations of ground-truth MaE samples in CAS(ME)2 are all within a normal range. In contrast, several MaE samples in SAMM abnormally exceed the general duration, making them practically difficult to spot during inference.
In summary, our proposed model can achieve a more advanced and balanced performance on both datasets, demonstrating the great generalization ability of our model. In addition, our method requires the detection of facial landmarks and the calculation of optical flows to form a head pose-irrelevant graph, making it more robust to variations such as random head angles. These merits demonstrate the potential of our method for future practical applications.